Instructions: Read each item below. Use the PVIF and FVIF tables and the simple interest formula...

  1. Instructions: Read each item below. Use the PVIF and FVIF tables and the simple interest formula to help answer the items below.
  1. Joel is going to put $2500 in a savings account at a local bank. The savings account will earn Joel 8% annually, simple interest. How much will Joel have in his account if he leaves the money there for 5 years? 10 years?
  2. If Joel had put the $2500 in an account in which the interest compounded annually, how would his account balance differ from the balances in #1 (At 5 years and 10 years)
  3. Cade has $5000 available to invest. He has found an opportunity to invest this money at 12% annually. If Cade left this money to grow at this rate (compounding) for 10 years, how much would he have at the end of the 10th year?
  4. If Cade were able to leave it for 15 years, what would the balance be?
  5. Cameron wants to have $5000 for a down payment on a vehicle. He plans to purchase the vehicle in approximately 4 years (at the time he thinks he’ll need to get rid of his current vehicle). If he can earn 9% interest, how much must he deposit now in order to have the money in 4 years?
  6. What would the present value of an account be if the future value is $20,000 and it has been compounding at a rate of 6% for the past 5 years?
  7. What should Jeff be willing to invest now in order to have $20,000 in 15 years if his investment will compound at a rate of 10% annually.
  8. Jordan wants to accumulate $1000 two years from now. If he can earn 6% on his investment, how much should he invest now?
  9. Maria would like to know the difference in account values if she:
    1. Invested $10,000 for 5 years @ 8%, simple interest
    2. Invested $10,000 for 5 years @ 8%, compounding interest
Answer #1

#1. Simple Interest = РxRx1/100

For 5 years

= 2500*8%*5 = $1000

Amount = Principal + Interest = 2500+1000 = $3500

For 10 years

= 2500*8%*10 = $2000

Amount = 2500+2000= $4500

#2. Amount at Compound Interest = P(1+R)T

For 5 years

= 2500(1+8%)5 = $3673.32

For 10 years

= 2500(1+8%)10 = $5397.312

#3. Amount at the end of 10th year = 5000(1+12%)10 = $15,529.241

#4. Amount at the end of 15th year = 5000(1+12%)15 = $27,367.83

#5. We will calculate the present value of the payment to be made after 4 years -

=> Present Value = Future Value / (1+i)n = $5000 / (1+9%)4 = $3542.13

He will have to deposit $3542.13 now to have $5000 at the end of 4th year compounded annually.

#6. PV = $20,000 / (1+6%)5 = 14,945.163

#7. PV = $20,000 / (1+10%)15 = $4787.84 must be invested by Jeff now to get $20,000 at the end of 15th year.

If he gets 6% interest , then PV = $20,000 / (1.06)15 = $8345.30

#8. a. Amount at Simple Interest = 10,000 + 10,000*8%*5 = $14,000

b. At compound interest = 10,000(1+8%)5 = $14,693.28

Difference in the account values = $693.28